# CARTOS: A Charging-Aware Real-Time Operating System for Intermittent Batteryless Devices

### v2

@@ V-A Hardware Setup
Fig. 7 depicts the hardware setup for our experiments. The main processing board is an Ambiq Apollo4 Evaluation board [36] equipped with a Cortex-M4 Processor, 2MB internal non-volatile MRAM, and 2.75MB low power SRAM. Power is supplied by an energy harvesting board comprising a 30mF, 100mF, or 470mF capacitor (will be explained in Sec V-B), LTC3588 energy harvester, and a 2W solar panel. An energy controller board with an ultra-low-power MSP430 MCU is used to control the power supply to the processing board and record experimental results. We have also integrated and configured a camera within the system to facilitate applications related to image processing as well as an environmental sensor to measure temperature, humidity, total volatile compound (TVOC) for environmental sensing applications. In our setup, the power-on and power-off thresholds are 4.04V and 2.9V respectively, determined by the datasheet when generating 1.8V. The low-voltage threshold is set to 3.0V accordingly. The difference between the power-off and low-voltage thresholds is to ensure enough energy for JIT checkpointing. The main processing board measures the capacitor voltage at runtime using the MCU’s internal ADC.
@@
While we used an external MCU (MSP430) in the energy controller to facilitate power supply control and experimental data recording, it could be replaced by an ultra-low-power programmable RTC (e.g., 14nA with Ambiq AM0815) or similar device capable of setting a GPIO pin for a programmable duration. In fact, even with MSP430, the power consumption of the energy controller is rather negligible ($<$0.5mW in peak usage) compared to that of the main processing board (9-93mW as shown in Table II). Hence, we did not consider the power overhead of the energy controller.
@@ V-B Real System Experiments
In the following, we evaluate CARTOS against state-of-the-art IPD scheduling methods under various experimental scenarios. All experiments in this section consider only single-task chains since most previous work in this domain does not thoroughly support task chains. [^5] Recall that chains in our work refer to the sequence of long-running threads, not that of short atomic blocks considered in some prior work [11]. To maintain consistency and repeatability, we employ several distinct harvesting conditions throughout the experiments (explained later). Each experiment begins when the capacitor reaches the “power-on” threshold and runs for four hyperperiods (480 seconds) to minimize the impact of initial capacitor charge. We used a solar panel and an LED lamp placed at a fixed distance apart. Our measurements show that the solar panel generates 8mW or 15mW of power when the LED lamp is on low or high setting, respectively.
Table II lists the taskset used in our experiments. This taskset includes preemptible tasks from the MiBench benchmark suite [43], along with two non-preemptible tasks: Sensor reading and Camera capture. Following common practice in real-time embedded systems, we used the rate-monotonic (RM) priority assignment, which assigns higher priorities ($\pi_{i}$) to tasks with shorter periods. For each task, the worst-case execution time ($C_{i}$) and average power consumption ($W_{i}$) were measured offline by executing each task 1000 times to obtain robust measurements. We compare CARTOS against InK [20], Quickrecall [8], CatNap [9], and Rtag [13]. These methods collectively represent the techniques listed in Table I or have been shown to outperform the others. Hence, this provides a comprehensive comparison of the efficacy of CARTOS across various scheduling scenarios.
@@ V-B1 Scheduler performance comparison in different charging conditions
InK and Rtag only consider non-preemptive scheduling [20, 13]. InK and Quickrecall use a best-effort approach [20, 8], i.e., execute tasks as soon as the device’s voltage reaches the power-on threshold and continue until the voltage drops to the power-off threshold. Therefore, we considered the capacitor size so that in all of the methods, at least one job of each task can complete its execution in one power cycle, i.e., $\frac{1}{2}*\mathbb{C}\times(V^{2}_{on}-V^{2}_{off})>\max_{i}\{(W_{i}-W_{S})\times C_{i}\}$, where $\mathbb{C}$ represents the capacitance, $V_{on}$ and $V_{off}$ are the power-on and power-off voltage thresholds, $W_{i}$ is the average power consumption of task $i$, $W_{S}$ is the power harvesting rate of the system, and $C_{i}$ is the worst-case execution time of task $i$. With this, the performance metric we considered is the ratio of successfully executed jobs to total jobs released since uncompleted jobs cannot produce correct output in IPDs.
@@
Fig. 8(b) demonstrates the scheduling behavior in moderate harvesting mode. In this scenario, only CARTOS successfully schedules all jobs, while the other methods result in missed deadlines, even for higher-priority tasks. InK lacks JIT checkpointing support and executes all tasks atomically. Therefore, when the voltage drops below the turn-off threshold, any running task under InK loses progress and must restart in the next power cycle, thereby wasting energy. Quickrecall uses a “best-effort” approach to task execution, only stopping job execution and performing JIT checkpointing at the low-voltage threshold. As a result, it fails to complete some non-preemptible jobs, forcing them to restart in the next power cycle and causing energy waste. CatNap supports JIT checkpointing for computational tasks and monitors charging requirements for peripheral tasks (‘events’ in [9]), allowing them to complete atomically without power interruption. However, it takes a hierarchical scheduling approach that strictly prioritizes peripheral tasks over computational tasks, nullifying the actual priority assignment and causing energy starvation for high-priority tasks. Rtag restricts all tasks to execute atomically, causing long blocking times for higher-priority tasks. It also faces an out-of-order energy consumption issue similar to CatNap, where lower-priority tasks consume energy that could otherwise be used by higher-priority tasks.
@@ V-B3 Scheduler and JIT Checkpointing Overhead of CARTOS
During the experiment, we measured the overhead caused by JIT checkpointing mechanism for preemptible tasks. The JIT checkpointing overhead was measured to be less than 0.088% in worst case scenario. It should be noted that the maximum JIT checkpointing overhead occurs when the smallest capacitor and the lowest charging rate were used, causing the largest number of power cycles. For the taskset used in previous experiments, the scheduler overhead was very negligible and could not be precisely measured by our equipment; however, it was estimated to be less than 1ms ($<$ 0.001%) by comparing device up time and task execution time. This shows that CARTOS delivers high-performance task scheduling on IPDs with minimal impact on system resources. We also measured the power consumption overhead associated with writing to and reading from MRAM compared to SRAM. To do so, we conducted 64MB of read/write operation and measured the average power consumption. Table III provides the overhead measurements in details.
@@
JIT: storing TCB and other states to MRAM 1.11 ms
JIT: storing taskset’s data to MRAM 1.46 ms
JIT: restoring all states and data from MRAM upon reboot 0.13 ms
@@ V-C Analytical Experiments
In this section, we compare the schedulability of our scheduler with Rtag [13], which performed the best among other methods in the previous section. We exclude other related works from this comparison because they either lack schedulability analysis, rely on overly simplified real-time analysis equations (e.g., considering only taskset utilization), and reportedly perform worse than Rtag [13].
@@
To determine the execution time of each task for our experiments, we choose the period randomly from 1s to 60s, find the task’s utilization using the UUniFast method [44], and then multiply this utilization by the task’s period and round the result to the nearest positive value with one decimal point accuracy. In simpler terms, we determine the task execution time as $C_{i}=\max(\lfloor 10\cdot T_{i}\cdot U_{i}\rceil/10,0.1)$.
@@ VI Conclusion
In this paper, we introduced CARTOS, a charging-aware real-time operating system for intermittently-powered betteryless devices. CARTOS leverages a mixed-preemption scheduling model to facilitate correct and efficient execution of periodic sensing applications, comprised of a chain of computational and peripheral tasks. For computational tasks, it schedules non-preemptively while ensuring forward progress through just-in-time (JIT) checkpointing. For peripheral tasks, it guarantees atomicity by scheduling them non-preemptively and charging sufficient energy to complete their jobs before execution. Furthermore, CARTOS includes energy management that computes the required charging time and utilizes a machine-learning-based energy predictor to update the harvesting rate to adapt to changing environmental conditions. We evaluated CARTOS through real system experiments as well as schedulability experiments. Our results showed that CARTOS outperfroms pevious methods by successfully scheduling periodic tasks regardless of the capacitor size and is able to schedule all higher priority jobs even in low charging situations with predictability. Furthermore, CARTOS outperformed state-of-the-art by up to 20% in schedulability across various experimental settings. For future work, we plan to expand the capabilities of CARTOS to accommodate a more diverse set of applications. This encompasses task synchronization and locking primitives, integrating with multi-core microcontrollers, and managing multiple energy buffers, all within the context of real-time operating systems and batteryless devices.

### v3

@@ V-A Hardware Setup
[…]
@@
While we used an external MCU (MSP430) in the energy controller to facilitate power supply control and experimental data recording, it could be replaced by an ultra-low-power programmable RTC (e.g., 14nA with Ambiq AM0815) or similar device capable of setting a GPIO pin for a programmable duration. In fact, even with MSP430, the power consumption of the energy controller is rather negligible (<<0.5mW in peak usage) compared to that of the main processing board (9-93mW as shown in Table II). Hence, we did not consider the power overhead of the energy controller.
@@ V-B Real System Experiments
In the following, we evaluate CARTOS against state-of-the-art IPD scheduling methods under various experimental scenarios. All experiments in this section consider only single-task chains since most previous work in this domain does not thoroughly support task chains. [^6] Recall that chains in our work refer to the sequence of long-running threads, not that of short atomic blocks considered in some prior work [colin2016chain]. To maintain consistency and repeatability, we employ several distinct harvesting conditions throughout the experiments (explained later). Each experiment begins when the capacitor reaches the “power-on” threshold and runs for four hyperperiods (480 seconds) to minimize the impact of initial capacitor charge. We used a solar panel and an LED lamp placed at a fixed distance apart. Our measurements show that the solar panel generates 8mW or 15mW of power when the LED lamp is on low or high setting, respectively.
Table II lists the taskset used in our experiments. This taskset includes preemptible tasks from the MiBench benchmark suite [MiBench_Benchmark], along with two non-preemptible tasks: Sensor reading and Camera capture. Following common practice in real-time embedded systems, we used the rate-monotonic (RM) priority assignment, which assigns higher priorities ($\pi_{i}$) to tasks with shorter periods. For each task, the worst-case execution time ($C_{i}$) and average power consumption ($W_{i}$) were measured offline by executing each task 1000 times to obtain robust measurements. We compare CARTOS against InK [Ink2018], Quickrecall [QuickRecall2014], CatNap [CatNap], and Rtag [MohsenIOTJ2021]. These methods collectively represent the techniques listed in Table I or have been shown to outperform the others. Hence, this provides a comprehensive comparison of the efficacy of CARTOS across various scheduling scenarios.
@@ V-B1 Scheduler performance comparison in different charging conditions
InK and Rtag only consider non-preemptive scheduling [Ink2018, MohsenIOTJ2021]. InK and Quickrecall use a best-effort approach [Ink2018, QuickRecall2014], i.e., execute tasks as soon as the device’s voltage reaches the power-on threshold and continue until the voltage drops to the power-off threshold. Therefore, we considered the capacitor size so that in all of the methods, at least one job of each task can complete its execution in one power cycle, i.e., $\frac{1}{2}*\mathbb{C}\times(V^{2}_{on}-V^{2}_{off})>\max_{i}\{(W_{i}-W_{S})\times C_{i}\}$, where $\mathbb{C}$ represents the capacitance, $V_{on}$ and $V_{off}$ are the power-on and power-off voltage thresholds, $W_{i}$ is the average power consumption of task $i$, $W_{S}$ is the power harvesting rate of the system, and $C_{i}$ is the worst-case execution time of task $i$. With this, the performance metric we considered is the ratio of successfully executed jobs to total jobs released since uncompleted jobs cannot produce correct output in IPDs.
@@
Fig. 8(b) demonstrates the scheduling behavior in moderate harvesting mode. In this scenario, only CARTOS successfully schedules all jobs, while the other methods result in missed deadlines, even for higher-priority tasks. InK lacks JIT checkpointing support and executes all tasks atomically. Therefore, when the voltage drops below the turn-off threshold, any running task under InK loses progress and must restart in the next power cycle, thereby wasting energy. Quickrecall uses a “best-effort” approach to task execution, only stopping job execution and performing JIT checkpointing at the low-voltage threshold. As a result, it fails to complete some non-preemptible jobs, forcing them to restart in the next power cycle and causing energy waste. CatNap supports JIT checkpointing for computational tasks and monitors charging requirements for peripheral tasks (‘events’ in [CatNap]), allowing them to complete atomically without power interruption. However, it takes a hierarchical scheduling approach that strictly prioritizes peripheral tasks over computational tasks, nullifying the actual priority assignment and causing energy starvation for high-priority tasks. Rtag restricts all tasks to execute atomically, causing long blocking times for higher-priority tasks. It also faces an out-of-order energy consumption issue similar to CatNap, where lower-priority tasks consume energy that could otherwise be used by higher-priority tasks.
@@ V-B3 Scheduler and JIT Checkpointing Overhead of CARTOS
During the experiment, we measured the overhead caused by JIT checkpointing mechanism for preemptible tasks. The JIT checkpointing overhead was measured to be less than 0.088% in worst case scenario. It should be noted that the maximum JIT checkpointing overhead occurs when the smallest capacitor and the lowest charging rate were used, causing the largest number of power cycles. For the taskset used in previous experiments, the scheduler overhead was very negligible and could not be precisely measured by our equipment. Instead, by comparing device uptime and task execution time, we could estimate the scheduler overhead (including voltage monitoring costs) to be less than 1ms (<< 0.001%). This shows that CARTOS delivers high-performance task scheduling on IPDs with minimal impact on system resources. We also measured the power consumption overhead associated with writing to and reading from MRAM compared to SRAM. To do so, we conducted 64MB of read/write operation and measured the average power consumption. Table III provides the overhead measurements in details.
@@
JIT: storing TCB and other states to MRAM 1.11 ms (0.104 mJ)
JIT: storing taskset’s data to MRAM 1.46 ms (0.137 mJ)
JIT: restoring states & data from MRAM on reboot 0.13 ms (0.013 mJ)
@@ V-C Analytical Experiments
In this section, we compare the schedulability of our scheduler with Rtag [MohsenIOTJ2021], which performed the best among other methods in the previous section. We exclude other related works from this comparison because they either lack schedulability analysis, rely on overly simplified real-time analysis equations (e.g., considering only taskset utilization), and reportedly perform worse than Rtag [MohsenIOTJ2021].
@@
To determine the execution time of each task for our experiments, we choose the period randomly from 1s to 60s, find the task’s utilization using the UUniFast method [UUnifast], and then multiply this utilization by the task’s period and round the result to the nearest positive value with one decimal point accuracy. In simpler terms, we determine the task execution time as $C_{i}=\max(\lfloor 10\cdot T_{i}\cdot U_{i}\rceil/10,0.1)$.
@@ VI Conclusion
In this paper, we introduced CARTOS, a charging-aware real-time operating system for intermittently-powered betteryless devices. CARTOS leverages a mixed-preemption scheduling model to facilitate correct and efficient execution of periodic sensing applications, comprised of a chain of computational and peripheral tasks. For computational tasks, it schedules non-preemptively while ensuring forward progress through just-in-time (JIT) checkpointing. For peripheral tasks, it guarantees atomicity by scheduling them non-preemptively and charging sufficient energy to complete their jobs before execution. Furthermore, CARTOS includes energy management that computes the required charging time and utilizes a machine-learning-based energy predictor to update the harvesting rate to adapt to changing environmental conditions.
We evaluated CARTOS through real system experiments as well as schedulability experiments. Our results showed that CARTOS outperforms previous methods by successfully scheduling periodic tasks regardless of the capacitor size and is able to schedule all higher priority jobs even in low charging situations with predictability. Furthermore, CARTOS outperformed state-of-the-art by up to 20% in schedulability across various experimental settings.
For future work, we plan to expand the capabilities of CARTOS to accommodate a more diverse set of applications. This encompasses task synchronization and locking primitives, integrating with multi-core microcontrollers, and managing multiple energy buffers, all within the context of real-time operating systems and batteryless devices.
[Uncaptioned image] Mohsen Karimi (Member, IEEE) received the Ph.D. degree in electrical and computer engineering from the University of California at Riverside, Riverside, CA, USA, in 2023. He is currently a Project Scientist with the Department of Electrical and Computer Engineering, University of California at Riverside, Riverside, CA, USA. His research interests include real-time task scheduling, embedded systems, cyber-physical systems, and machine learning, with a focus on designing energy harvesting devices, real-time task scheduling, resource management, and machine learning techniques for batteryless devices.
[Uncaptioned image] Yidi Wang (Member, IEEE) Yidi Wang is currently an Assistant Professor in the Department of Computer Science and Engineering at Santa Clara University. Before joining SCU in 2024, she was a Postdoc in the Department of Electrical and Computer Engineering at the University of California, Riverside. She earned her Ph.D. in Electrical Engineering from UC Riverside in June 2023. Her primary research interests lie in real-time embedded and cyber-physical systems with a particular focus on real-time scheduling on heterogeneous systems in resource-constrained environments.
[Uncaptioned image] Youngbin Kim (Member, IEEE) is a Senior Researcher in Electronics and Telecommunications Research Institute, Daejeon, South Korea. He received his B.S. degree in 2012 and Ph.D. degree in 2021, both from Yonsei University, Seoul, South Korea. His research is in the area of embedded systems, with a special interest in intermittent computing, neural network acceleration, and reliability.
[Uncaptioned image] Yoojin Lim (Member, IEEE) received the B.S. degree in computer science and computer engineering and the M.S. degree in information technology from Handong Global University, South Korea, in 2004, and the Ph.D. degree in business information technology from Kookmin University, South Korea, in 2016. From 2004 to 2008, he was a Senior Software Engineer with LG Electronics, South Korea. Since 2017, he has been a postdoctoral researcher and senior researcher with Electronics and Telecommunications Research Institute, Daejeon, South Korea, and he is currently a project leader of autonomous energy-driven computing system software research. His research interests include energy-driven intermittent computing, embedded software, hypervisors, cyber-physical systems, and mobility systems.
[Uncaptioned image] Hyoseung Kim (Member, IEEE) received the Ph.D. degree in electrical and computer engineering from Carnegie Mellon University, Pittsburgh, PA, USA, in 2016. He is an Associate Professor with the Department of Electrical and Computer Engineering, University of California at Riverside, Riverside, CA, USA. His research interests are in real-time embedded and cyber-physical systems, especially at the intersection of systems software, parallel computing platforms, and analytical techniques. Dr. Kim is a recipient of the NSF CAREER Award, the Fulbright Scholarship Award, and the Best Paper Awards from RTAS and RTCSA. For more information, please visit http://www.ece.ucr.edu/~hyoseung/.